# Diffusion Models: A Comprehensive Survey of Methods and Applications

### v13

@@ 7.4.3. Text-to-3D Generation
3D content generation (Poole et al., 2022; Lin et al., 2022; Xu et al., 2022a; Jun and Nichol, 2023) has been in high demand for a wide range of applications, including gaming, entertainment, and robotics simulation. Augmenting 3D content generation with natural language could considerably help with both novices and experienced artists. DreamFusion (Poole et al., 2022) adopts a pre-trained 2D text-to-image diffusion model to perform text-to-3D synthesis. It optimizes a randomly-initialized 3D model (a Neural Radiance Field, or NeRF) with a probability density distillation loss, which utilizes a 2D diffusion model as a prior for optimization of a parametric image generator. To obtain fast and high-resolution optimization of NeRF, Magic3D (Lin et al., 2022) proposes a two-stage diffusion framework built on cascaded low-resolution image diffusion prior and high-resolution latent diffusion prior. In order to achieve high-fidelity 3D creation, Make-It-3D (Tang et al., 2023) optimizes a neural radiance field by incorporating constraints from the reference image at the frontal view and diffusion prior at novel views, enhancing the coarse model into textured point clouds and increasing realism with diffusion prior and high-quality textures from the reference image. ProlificDreamer (Wang et al., 2023) presents Variational Score Distillation (VSD), optimizing a distribution of 3D scenes based on textual prompts as random variables to closely align the distribution of rendered images from all perspectives with a pretrained 2D diffusion model, using KL divergence as the measure. As demonstrated in Fig. 14, IPDreamer further proposes a novel 3D object synthesis framework that enables users to create controllable and high-quality 3D objects effortlessly. It excels in synthesizing a high-quality 3D object which can greatly align with a provided complex image prompt.
@@ 7.4.5. Text-to-Video Generation
Tremendous recent progress in text-to-image diffusion-based generation motivates the development of text-to-video generation (Singer et al., 2022; Ho et al., 2022a; Wu et al., 2022a). Make-A-Video (Singer et al., 2022) proposes to extend a diffusion-based text-to-image model to text-to-video through a spatiotemporally factorized diffusion model. It leverages joint text-image prior to bypass the need for paired text-video data, and further presents super-resolution strategies for high-definition, high frame-rate text-to-video generation. Imagen Video (Ho et al., 2022a) generates high definition videos by designing a cascaded video diffusion models, and transfers some findings that work well in the text-to-image setting to video generation, including frozen T5 text encoder and classifier-free guidance. Tune-A-Video (Wu et al., 2022a) introduces one-shot video tuning for text-to-video generation, which eliminates the burden of training with large-scale video datasets. It employs efficient attention tuning and structural inversion to significantly enhance temporal consistency. Text2Video-Zero (Khachatryan et al., 2023) achieves zero-shot text-to-video synthesis using a pretrained text-to-image diffusion model, ensuring temporal consistency through motion dynamics in latent codes and cross-frame attention. Its goal is to enable affordable text-guided video generation and editing without additional fine-tuning. FateZero (Qi et al., 2023) is the first framework for temporal-consistent zero-shot text-to-video editing using pre-trained text-to-image diffusion model. It fuses the attention maps in the DDIM inversion and generation processes to maximally preserve the consistency of motion and structure during editing. ContextDiff (Yang et al., 2024f) incorporates the cross-modal context information about the interactions between text condition and video sample into forward and reverse processes, forming a forward-backward consistent video diffusion model for text-to-video generation.
@@ 7.5.2. Time Series Forecasting
Time series forecasting is the task of forecasting or predicting the future value over a period of time. Neural methods have recently become widely-used for solving the prediction problem with univariate point forecasting methods (Oreshkin et al., 2019) or univariate probabilistic methods (Salinas et al., 2019). In the multivariate setting, we also have point forecasting methods (Li et al., 2019a) as well as probabilistic methods, which explicitly model the data distribution using Gaussian copulas (Salinas et al., 2020), GANs (Yoon et al., 2019), or normalizing flows (Rasul et al., 2020). TimeGrad (Rasul et al., 2021b) presents an autoregressive model for forecasting multivariate probabilistic time series, which samples from the data distribution at each time step through estimating its gradient. It utilizes diffusion probabilistic models, which are closely connected with score matching and energy-based methods. Specifically, it learns gradients by optimizing a variational bound on the data likelihood and transforms white noise into a sample of the distribution of interest through a Markov chain using Langevin sampling (Song and Ermon, 2019) during inference time.
@@ 8. Future Directions
From Stable Diffusion to ChatGPT, Artificial Intelligence Generated Content (AIGC) has gained much attention in both academic and industrial circles. Generative Pre-Training is the core technique in GPT-1/2/3/4 (Radford et al., 2018, 2019; Ouyang et al., 2022; OpenAI, 2023a) and (Visual) ChatGPT (Wu et al., 2023), which exhibits promising generation performance and surprising emergent abilities (Wei et al., 2022) equipped with Large Language Models (LLMs) (Touvron et al., 2023) and Visual Foundation Models (Bommasani et al., 2021; Yu et al., 2022b; Yuan et al., 2021). It is interesting to transfer the generative pre-training (decoder-only) from GPT series to diffusion model class, evaluate the diffusion-based generation performance at scale, and analyse the emergent abilities of diffusion foundation models. Furthermore, combining LLMs with diffusion models have been proved to be a new promising direction (Yang et al., 2024c, e).

### v14

@@ 7.4.3. Text-to-3D Generation
3D content generation (Poole et al., 2022; Lin et al., 2022; Xu et al., 2022a; Jun and Nichol, 2023) has been in high demand for a wide range of applications, including gaming, entertainment, and robotics simulation. Augmenting 3D content generation with natural language could considerably help with both novices and experienced artists. DreamFusion (Poole et al., 2022) adopts a pre-trained 2D text-to-image diffusion model to perform text-to-3D synthesis. It optimizes a randomly-initialized 3D model (a Neural Radiance Field, or NeRF) with a probability density distillation loss, which utilizes a 2D diffusion model as a prior for optimization of a parametric image generator. To obtain fast and high-resolution optimization of NeRF, Magic3D (Lin et al., 2022) proposes a two-stage diffusion framework built on cascaded low-resolution image diffusion prior and high-resolution latent diffusion prior. In order to achieve high-fidelity 3D creation, Make-It-3D (Tang et al., 2023) optimizes a neural radiance field by incorporating constraints from the reference image at the frontal view and diffusion prior at novel views, enhancing the coarse model into textured point clouds and increasing realism with diffusion prior and high-quality textures from the reference image. ProlificDreamer (Wang et al., 2023) presents Variational Score Distillation (VSD), optimizing a distribution of 3D scenes based on textual prompts as random variables to closely align the distribution of rendered images from all perspectives with a pretrained 2D diffusion model, using KL divergence as the measure. IPDreamer (Zeng et al., 2023) further proposes a novel 3D object synthesis framework that enables users to create controllable and high-quality 3D objects effortlessly. It excels in synthesizing a high-quality 3D object which can greatly align with a provided complex image prompt.
Modeling compositional 3D data distribution is a fundamental and critical task for generative models. Current feed-forward methods (Shue et al., 2023; Shi et al., 2024) are primarily capable of generating single objects and face challenges when creating more complex scenes containing multiple objects due to limited training data. Recently, a series of learnable-layout compositional methods have been proposed (Epstein et al., 2024; Vilesov et al., 2023; Han et al., 2024; Chen et al., 2024; Gao et al., 2024) . These methods combine multiple object-ad-hoc radiance fields and then optimize the positions of the radiance fields from external feedback. For example, Epstein et al. (2024) propose learning a distribution of reasonable layouts based solely on the knowledge from a large pre-trained text-to-image model. Vilesov et al. (2023) introduce an optimization method based on Monte-Carlo sampling and physical constraints. However, these forms of layout guidance are relatively coarse and not expressive enough for fine-grained control. Yang et al. (2024) (Yang et al., 2024g) address this problem by incorporating semantic embeddings that ensure view consistency and distinctly differentiate objects into SDS processes (namely SemanticSDS (Yang et al., 2024g)), which are flexible and expressive for optimizing 3D scenes. As shown in Fig. 15, SemanticSDS (Yang et al., 2024g) can achieve superior precision and quality in compositional text-to-3D generation over existing methods.
@@ 7.4.5. Text-to-Video Generation
Tremendous recent progress in text-to-image diffusion-based generation motivates the development of text-to-video generation (Singer et al., 2022; Ho et al., 2022a; Wu et al., 2022a). Make-A-Video (Singer et al., 2022) proposes to extend a diffusion-based text-to-image model to text-to-video through a spatiotemporally factorized diffusion model. It leverages joint text-image prior to bypass the need for paired text-video data, and further presents super-resolution strategies for high-definition, high frame-rate text-to-video generation. Imagen Video (Ho et al., 2022a) generates high definition videos by designing a cascaded video diffusion models, and transfers some findings that work well in the text-to-image setting to video generation, including frozen T5 text encoder and classifier-free guidance. Tune-A-Video (Wu et al., 2022a) introduces one-shot video tuning for text-to-video generation, which eliminates the burden of training with large-scale video datasets. It employs efficient attention tuning and structural inversion to significantly enhance temporal consistency. Text2Video-Zero (Khachatryan et al., 2023) achieves zero-shot text-to-video synthesis using a pretrained text-to-image diffusion model, ensuring temporal consistency through motion dynamics in latent codes and cross-frame attention. Its goal is to enable affordable text-guided video generation and editing without additional fine-tuning. FateZero (Qi et al., 2023) is the first framework for temporal-consistent zero-shot text-to-video editing using pre-trained text-to-image diffusion model. It fuses the attention maps in the DDIM inversion and generation processes to maximally preserve the consistency of motion and structure during editing. ContextDiff (Yang et al., 2024h) incorporates the cross-modal context information about the interactions between text condition and video sample into forward and reverse processes, forming a forward-backward consistent video diffusion model for text-to-video generation.
@@ 7.5.2. Time Series Forecasting
Time series forecasting is the task of forecasting or predicting the future value over a period of time. Neural methods have recently become widely-used for solving the prediction problem with univariate point forecasting methods (Oreshkin et al., 2019) or univariate probabilistic methods (Salinas et al., 2019). In the multivariate setting, we also have point forecasting methods (Li et al., 2019a) as well as probabilistic methods, which explicitly model the data distribution using Gaussian copulas (Salinas et al., 2020), GANs (Yoon et al., 2019), or normalizing flows (Rasul et al., 2020). TimeGrad (Rasul et al., 2021b) presents an autoregressive model for forecasting multivariate probabilistic time series, which samples from the data distribution at each time step through estimating its gradient. It utilizes diffusion probabilistic models, which are closely connected with score matching and energy-based methods. Specifically, it learns gradients by optimizing a variational bound on the data likelihood and transforms white noise into a sample of the distribution of interest through a Markov chain using Langevin sampling (Song and Ermon, 2019) during inference time. To handle complex time series forecasting, as illustrated in Fig. 17, Liu et al. (2024) for the first time introduce Retrieval- Augmented Time series Diffusion (RATD) (Liu et al., 2024), allowing for greater utilization of the dataset and providing meaningful guidance in the denoising process.
@@ 8. Future Directions
From Stable Diffusion to ChatGPT, Artificial Intelligence Generated Content (AIGC) has gained much attention in both academic and industrial circles. Generative Pre-Training is the core technique in GPT-1/2/3/4 (Radford et al., 2018, 2019; Ouyang et al., 2022; OpenAI, 2023a) and (Visual) ChatGPT (Wu et al., 2023), which exhibits promising generation performance and surprising emergent abilities (Wei et al., 2022) equipped with Large Language Models (LLMs) (Touvron et al., 2023) and Visual Foundation Models (Bommasani et al., 2021; Yu et al., 2022b; Yuan et al., 2021). It is interesting to transfer the generative pre-training (decoder-only) from GPT series to diffusion model class, evaluate the diffusion-based generation performance at scale, and analyse the emergent abilities of diffusion foundation models. Furthermore, combining LLMs with diffusion models have been proved to be a new promising direction (Yang et al., 2024d, f).